OpenStudy (anonymous):

Let f(x) = x^2 + x + 2 and g(x) = 2x^2 + 5. Find f(g(x)). Show each step of your work.

OpenStudy (anonymous):

@Vocaloid I just need to learn how to do the problem, not the answer.

Nnesha (nnesha):

replace x in f(x) function by g(x)

OpenStudy (anonymous):

I did that, I got: (2x^2+5)^2+2x^2+2

Nnesha (nnesha):

you forgot something \[ \huge\rm (2x^2+5)^2+\color{blue}{x}+2\] x = 2x^2 + 5

Nnesha (nnesha):

you forgot 5

OpenStudy (anonymous):

did I? Sorry.

OpenStudy (anonymous):

So to solve, is it 2x^4 or 4x^4

Nnesha (nnesha):

\[ \huge\rm (2x^2+5)^2+\color{blue}{2x^2 +5}+2\]

Nnesha (nnesha):

\[ \huge\rm \color{reD}{x^2}+\color{blue}{x}+2\] \[ \huge\rm \color{ReD}{(2x^2+5)^2}+\color{blue}{2x^2 +5}+2\] x is 2x^2 + 5

OpenStudy (anonymous):

Yes, so now we have to square x.

Nnesha (nnesha):

yes( 2x^2 +5)^2 is same as (2x^2 +5) (2x^2 +5) so foil it

OpenStudy (anonymous):

\[4x^4+20x^2+25\]

OpenStudy (anonymous):

right?

Nnesha (nnesha):

yes that's right

OpenStudy (anonymous):

So now we combine like terms?

Nnesha (nnesha):

yesp

OpenStudy (anonymous):

\[4x^4+20x^2+25+2x^2+5+2\]

OpenStudy (anonymous):

turns into

OpenStudy (anonymous):

\[4x^4+22x^2+32\]

OpenStudy (anonymous):

Right?

Nnesha (nnesha):

yep

OpenStudy (anonymous):

So now I need to factor?

Nnesha (nnesha):

nope that's ur answer

OpenStudy (anonymous):

Oh, is it? Good, my problem was I tried factoring that and had no idea how to do it.

OpenStudy (anonymous):

Thanks! :D

Nnesha (nnesha):

my pleasure and yes that's it they wants us to find f(g(x)) not factoring

OpenStudy (anonymous):

I fanned, so if I need anything, I'll make sure to tag you.

Nnesha (nnesha):

thanks i'll try my best :=)

OpenStudy (anonymous):

@Nnesha I have one more problem that seems pretty easy.

OpenStudy (anonymous):

I think I may know how to solve it but graphing and radicals are my weakpoint.

OpenStudy (anonymous):

Describe how to transform \[(\sqrt[6]{x ^{5}})^{7}\] into an expression with a rational exponent

Nnesha (nnesha):

alright remember this exponent rule \[\huge\rm \sqrt[m]{x^n} = x^\frac{ n }{ m }\] you can convert root to exponent form

OpenStudy (anonymous):

Yes I know this, it would be \[(x^\frac{ 5 }{ 6 })^7\]

OpenStudy (anonymous):

then you multiply 7 and 5?

Nnesha (nnesha):

yep! right you already know

Nnesha (nnesha):

\[\huge\rm (x^m)^n = x^{ m \times n}\]

OpenStudy (anonymous):

\[x^\frac{ 35 }{ 6 }\]

OpenStudy (anonymous):

thats a 35/6 btw

Nnesha (nnesha):

yep

OpenStudy (anonymous):

And that is my answer?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

K, thanks. Just wanted to check my answer.

Nnesha (nnesha):

your answer is correct you're a mathematician :=)